# A Concept-Based Explainability Framework for Large Multimodal Models

### v1

@@ 5 Conclusion
Appendix A Further implementation details
A.1 Dictionary learning details
@@
Table 3: Number of samples training/testing samples for each token
@@
A.2 CLIPScore/BERTScore evaluation
@@
We found that computing BERTScores from comma-separated words and captions is unreliable. Instead, we adopted a method using the LLaMA-3-8B instruct model to construct coherent phrases from a set of grounded words, $\mathbf{T}_{k}$. Specifically, we provide the LLaMA model with instructions to describe a scene using a designated set of words, for which we also supply potential answers. This instruction is similarly applied to another set of words, but without providing answers. The responses generated by LLaMA are then compared to the captions $y$ using BERTScore. The instruction phrase and an example of the output are detailed in 4. The highest matching score between the generated phrases and the captions of a test sample determines the score assigned to the concept $u_{k}$. This approach ensures that the evaluation accurately reflects coherent and contextually integrated language use. The metric calculation from embeddings follows the established guidelines outlined in [47]. Our adapted implementation is based on BERTScore official repository, and we use the default Roberta-large model to generate embeddings.
@@
A.3 Resources
Compute usage
Each experiment to analyze a token with a selected dictionary learning method is conduced on a single RTX5000 (24GB)/ RTX6000 (48GB)/ TITAN-RTX (24GB) GPU. Within dictionary learning, generating visualizations and projecting test data, the majority of time is spent in loading the data/models and extracting the representations. For analysis of a single token with  3000 training samples, it takes around 10-15 mins for this complete process. Evaluation for CLIPScore/BERTScore ar also conducted using the same resources. Evaluating CLIPScore for 500 (image, grounded-words) pairs takes around 5 mins.
@@ A.3 Resources
Appendix B Quantitative evaluation for more tokens
We provide test data mean CLIPScore and BERTScore for top-1 activating concept for all baselines and more tokens: Baby, Car, Cake, and Bear in 5 (results in the main paper are reported for tokens Dog, Bus, Train, Cat in the 1). We observe that we consistently obtain higher scores across our approach. We also employ other metrics such as overlap between grounded words to illustrate the superiority of our method over the simple baseline; this metric is reported in 6. As previously noted, we observe a high overlap between grounded words with KMeans and Simple baselines compared to Semi-NMF/PCA. A low overlap should be encouraged, as it indicates the discovery of diverse and disentangled concepts in the dictionary.
@@ Appendix B Quantitative evaluation for more tokens
Table 5: Test data mean CLIPScore and BERTScore for top-1 activating concept CLIPScore denoted as CS, BERTScore denoted as BS for all concept extraction baselines considered. Analysis for layer $L=31$. Best score indicated in bold and second best is underlined.
@@
Table 6: Overlap/entanglement between grounded words of learnt concepts for different dictionary learning methods. Results are for additional tokens. Lower is better. Best score indicated in bold and second best is underlined.
B.1 Statistical significance
The statistical significance of Semi-NMF w.r.t all other baselines and variants, for CLIPScore/BERTScore evaluation to understand representations of test samples is given in Tab/ 7 (for all tokens separately). We report the $p$-values for an independent two sided T-test with null hypothesis that mean performance is the same between Semi-NMF and the respective system. The results for Semi-NMF are almost always significant compared to Rnd-Words, Noise-Imgs, PCA. However for these metrics, Simple baseline, K-Means and Semi-NMF all perform competitively and better than other systems. Within these three systems the significance depends on the target token, but are often not significant in many cases.
@@ B.1 Statistical significance
Appendix C Additional visualizations
C.1 Concept grounding
The visual/textual grounding for all tokens in Tab. 1 are given in Figs. 7 (‘Dog’), 8 (‘Cat’), 9 (‘Bus’), 10 (‘Train’). All the results extract $K=20$ concepts from layer $L=31$. Similar to our analysis for token ‘Dog’ in main paper, for a variety of target tokens our method extracts diverse and multimodally coherent concepts encoding various aspects related to the token.
@@ C.1 Concept grounding
C.2 Local interpretations
Here, we qualitatively analyze the local interpretations of various decomposition methods, including PCA, k-means, semi-NMF, and the simple baseline strategy. We select these four as they produce coherent grounding compared to Rnd-Words and Noise-Img baselines. We decompose test sample representations on our learnt dictionary and visualize the top three activating components. Note that in the case of KMeans and Simple baseline, the projection maps a given test representation to a single element of the concept dictionary, the one closest to it. However, for uniformity we show the three most closest concept vectors for both. Figs. 11, 12, 13, 14, 15 are dedicated to interpretations of a single sample each, for all four concept extraction methods.
@@ C.2 Local interpretations
Appendix D Qualitative analysis for different layers
@@
Appendix E Analysis for visual tokens
@@ Appendix F Limitations
Appendix G Broader societal impact

### v3

@@ 5 Conclusion
Acknowledgments and Disclosure of Funding
We would like to thank Thomas Fel for his valuable comments on the paper. This work has been partially supported by ANR grant VISA DEEP (ANR-20-CHIA-0022), and HPC resources of IDRIS under the file number AD011014947, allocated by GENCI.
Appendix A Experiments on other LMMs
This section covers our experiments on other types of multimodal models. First, we test our approach on LLaVA to demonstrate that our approach generalizes to more recent networks that also fine-tune $f_{LM}$ on multimodal data. We also test our method on other variants of DePALM with non-CLIP visual encoders to observe their effect on CLIPScore.
A.1 Experiments with LLaVA
We conduct further experiments on LLaVA [28], a popular open-source LMM to demonstrate the generality of our method. The model uses a CLIP-ViT-L-336px visual encoder ($f_{V}$), a 2-layer linear connector ($C$) that outputs $N_{V}=576$ visual tokens, and a Vicuna-7B language model ($f_{LM}$, 32 layers). We use identical hyperparameters as for DePALM ($K=20,\lambda=1,L=31$). We report the test CLIPScore for top-1 activating concept, for Rnd-Words, Noise-Imgs, Simple and Semi-NMF, GT-captions in Tab. 3, and Overlap score for non-random baselines in Tab. 4. Quantitatively, we obtain consistent results to those observed for DePALM. Semi-NMF extracts most balanced concept dictionaries with high multimodal correspondence and low overlap. Qualitatively too, the method functions consistently and is able to extract concepts with meaningful multimodal grounding.
Dog CS top-1 ($\uparrow$) 0.537 $\pm$ 0.03 0.530 $\pm$ 0.05 0.567 $\pm$ 0.08 0.595 $\pm$ 0.07 0.777 $\pm$ 0.06
BS top-1 ($\uparrow$) 0.205 $\pm$ 0.07 0.227 $\pm$ 0.06 0.331 $\pm$ 0.07 0.305 $\pm$ 0.07 0.519 $\pm$ 0.11
Bus CS top-1 ($\uparrow$) 0.509 $\pm$ 0.04 0.487 $\pm$ 0.05 0.619 $\pm$ 0.06 0.591 $\pm$ 0.08 0.742 $\pm$ 0.05
BS top-1 ($\uparrow$) 0.198 $\pm$ 0.07 0.253 $\pm$ 0.06 0.319 $\pm$ 0.04 0.306 $\pm$ 0.06 0.460 $\pm$ 0.10
Train CS top-1 ($\uparrow$) 0.518 $\pm$ 0.03 0.505 $\pm$ 0.04 0.633 $\pm$ 0.05 0.640 $\pm$ 0.07 0.725 $\pm$ 0.05
BS top-1 ($\uparrow$) 0.177 $\pm$ 0.07 0.221 $\pm$ 0.04 0.310 $\pm$ 0.05 0.293 $\pm$ 0.05 0.432 $\pm$ 0.08
Cat CS top-1 ($\uparrow$) 0.536 $\pm$ 0.03 0.545 $\pm$ 0.04 0.625 $\pm$ 0.06 0.621 $\pm$ 0.07 0.795 $\pm$ 0.05
BS top-1 ($\uparrow$) 0.142 $\pm$ 0.06 0.235 $\pm$ 0.05 0.306 $\pm$ 0.06 0.329 $\pm$ 0.07 0.540 $\pm$ 0.11
Table 3: Concept extraction on LLaVA-v1.5: Test data mean CLIPScore reported for top-1 activating concept for same baselines and tokens as in main paper table 1. Higher scores are better. Best score in bold, second best is underlined.
Token Simple PCA KMeans Semi-NMF
Dog 0.435 0.008 0.429 0.149
Bus 0.464 0.010 0.518 0.124
Train 0.315 0.024 0.382 0.087
Cat 0.479 0.013 0.554 0.166
Table 4: Overlap evaluation (LLaVA). Lower is better. Best score in bold, second best underlined.
Qualitative results and Saliency maps
We also show qualitative examples of concepts extracted for token ‘Dog’ in Fig. 7. More examples for other ‘Cat’ and ‘Train’ tokens are given in Fig. 9 and 10. Interestingly, since LLaVA uses a connector $C$ that contains two linear layers, the visual tokens as processed inside $f_{LM}$ preserve the notion of specific image patch representations, i.e. $N_{V}=576$ tokens denoting representations for 576 (24 $\times$ 24) input patches. This allows us to further explore a simple and computationally cheap strategy of generating saliency maps to highlight which regions a concept vector activates on. To do this one can simply compute the inner product of any given concept vector $u_{k}$ with all visual token representations from corresponding layer $L$, i.e. $u_{k}^{T}[h^{1}_{(L)},...,h^{N_{V}}_{(L)}]$. This can be upscaled to the input image size to visualize the saliency map. We illustrate some qualitative outputs on concepts from ‘Dog’ in Fig. 8. .
Figure 7: Multimodal grounding for example concepts for ’Dog’ token (layer 31) on LLaVA.
Figure 8: Examples of generating visual concept saliency maps for two ‘Dog’ concepts for LLaVA. Red denotes high activations, blue denotes low activation (bottom row)
Figure 9: Multimodal grounding for example concepts for ’Cat’ token (layer 31) on LLaVA.
Figure 10: Multimodal grounding for example concepts for ’Train’ token (layer 31) on LLaVA.
A.2 DePALM with ViT visual encoders
We further test LMMs which do not contain a CLIP visual encoder to confirm that the high CLIPScore is not due to use of CLIP visual encoders. To test this, we experiment on two different DePALM models with frozen visual encoders different from CLIP, a frozen ViT-L encoder trained on ImageNet [11] and another frozen ViT-L trained as a masked autoencoder (MAE) [19]. Both LMMs use the same pretrained OPT-6.7B language model. Collectively, the three encoders (including CLIP) are pretrained for three different types of objectives. We use Semi-NMF to extract concept dictionaries, with all hyperparameters identical. The results are reported in tables below. ’Rnd-Words’ and ’GT-captions’ references are reported for each LMM separately, although they are very close to the ones in main paper. The "ViT-L (CLIP)" baseline denotes our system from the main paper that uses CLIP encoder. Importantly, we still obtain similar test CLIPScores as with CLIP visual encoder. The concept dictionaries still possess meaningful multimodal grounding. Many concepts also tend to be similar as for CLIP visual encoder, further indicating that processing inside language model plays a major role in the discovery of multimodally grounded concepts.
Token Rnd-Words ViT-L (ImageNet) ViT-L (CLIP) GT-captions
Dog 0.514 $\pm$ 0.05 0.611 $\pm$ 0.09 0.610 $\pm$ 0.09 0.783 $\pm$ 0.06
Bus 0.498 $\pm$ 0.05 0.644 $\pm$ 0.07 0.634 $\pm$ 0.08 0.739 $\pm$ 0.05
Train 0.494 $\pm$ 0.05 0.617 $\pm$ 0.07 0.646 $\pm$ 0.07 0.728 $\pm$ 0.05
Cat 0.539 $\pm$ 0.05 0.628 $\pm$ 0.07 0.627 $\pm$ 0.06 0.794 $\pm$ 0.06
Table 5: Test CLIPScore evaluation for DePALM with ViT-L frozen image encoder trained on ImageNet: Scores reported for top-1 activating concept of Semi-NMF for Rnd-Words, GT-captions and ViT-L (CLIP) which denotes the system in main text.
Token Rnd-Words ViT-L (MAE) ViT-L (CLIP) GT-captions
Dog 0.515 $\pm$ 0.05 0.602 $\pm$ 0.07 0.610 $\pm$ 0.09 0.784 $\pm$ 0.06
Bus 0.501 $\pm$ 0.05 0.627 $\pm$ 0.07 0.634 $\pm$ 0.08 0.737 $\pm$ 0.05
Train 0.483 $\pm$ 0.06 0.618 $\pm$ 0.08 0.646 $\pm$ 0.07 0.726 $\pm$ 0.05
Cat 0.541 $\pm$ 0.04 0.629 $\pm$ 0.09 0.627 $\pm$ 0.06 0.795 $\pm$ 0.06
Table 6: Test CLIPScore evaluation for DePALM with ViT-L frozen image encoder trained as masked autoencoder (MAE): Scores reported for top-1 activating concept of Semi-NMF for Rnd-Words, GT-captions and ViT-L (CLIP) which denotes the system in main text.
Appendix B Analyzing polysemanticity in the learnt concepts
We conducted a preliminary qualitative study on some concept vectors in the dictionary learnt for token "Dog" (DePALM model), to analyze if these concept vectors tend to activate strongly for a specific semantic concept (monosemantic) or multiple semantic concepts (polysemantic). In particular, we first manually annotated the 160 test samples for "Dog" for four semantic concepts, for which we knew we had concept vectors in our dictionary, namely "Hot dog" (Concept 2, row 1, column 2 in Fig. 7), "Black dog" (Concept 20, row 10, column 2 in Fig. 7), "Brown/orange dog" (Concept 6, row 3, column 2 in Fig. 7), and "Bull dog" (Concept 15, row 8, column 1 in Fig. 7). For a given semantic concept, we call this set $C_{true}$. Then, for its corresponding concept vector $u_{k}$ we find the set of test samples for which $u_{k}$ activates greater than a threshold $\tau$. This threshold was set to half of its maximum activation over test samples. We call this set of samples $C_{top}$. To estimate specificity of the concept vector we compute how many samples in $C_{top}$ lie in the ground-truth set, i.e. $|C_{top}|\cap|C_{true}|/|C_{top}|$.
We found Concept 2 ("Hot dog") to be most monosemantic with 100% specificity. For Concept 20 ("Black dog") too, we found high specificity of 93.3%. For concept 15 ("Bull dog") we observed the lowest specificity of 50%. This concept also activated for test images with toy/stuffed dogs. Interestingly, the multimodal grounding of concept 15 already indicates this superposition with maximum activating samples also containing images of ’toy dogs’. Concept 6 ("Brown/orange dog") is somewhere in between, with 76% specificity. This concept vector also activated sometimes for dark colored dogs, which wasn’t apparent from its multimodal grounding.
Prominent or distinct semantic concepts seem to be captured by more specific/monosemantic concept vectors, while more sparsely present concepts seem at risk to be superposed resulting in a more polysemantic concept vector capturing them. It is also worth noting that the multimodal grounding can be a useful tool in some cases to identify polysemanticity in advance.
Appendix C Further implementation details
C.1 Dictionary learning details
@@
Table 7: Number of samples training/testing samples for each token for DePALM
@@
C.2 CLIPScore/BERTScore evaluation
@@
We found that computing BERTScores from comma-separated words and captions is unreliable. Instead, we adopted a method using the LLaMA-3-8B instruct model to construct coherent phrases from a set of grounded words, $\mathbf{T}_{k}$. Specifically, we provide the LLaMA model with instructions to describe a scene using a designated set of words, for which we also supply potential answers. This instruction is similarly applied to another set of words, but without providing answers. The responses generated by LLaMA are then compared to the captions $y$ using BERTScore. The instruction phrase and an example of the output are detailed in 8. The highest matching score between the generated phrases and the captions of a test sample determines the score assigned to the concept $u_{k}$. This approach ensures that the evaluation accurately reflects coherent and contextually integrated language use. The metric calculation from embeddings follows the established guidelines outlined in [50]. Our adapted implementation is based on BERTScore official repository, and we use the default Roberta-large model to generate embeddings.
@@
C.3 Resources
DePALM experiments compute
Each experiment to analyze a token with a selected dictionary learning method is conduced on a single RTX5000 (24GB)/ RTX6000 (48GB)/ TITAN-RTX (24GB) GPU. Within dictionary learning, generating visualizations and projecting test data, the majority of time is spent in loading the data/models and extracting the representations. For analysis of a single token with $\approx$ 3000 training samples, it takes around 10-15 mins for this complete process. Evaluation for CLIPScore/BERTScore are also conducted using the same resources. Evaluating CLIPScore for 500 (image, grounded-words) pairs takes around 5 mins. The BERTScore evaluation is in contrast more expensive, consuming around 150 mins for 500 pairs.
LLaVA experiments compute
Each experiment to extract a concept dictionary for LLaVA was conducted on a single A100 (80GB) GPU. Representation extraction process for LLaVA is more expensive compared to DePALM consuming around 90 mins for $\approx$ 3000 samples. The remaining aspects of dictionary learning, multimodal grounding, representation projection etc. remains relatively cheap. The CLIPScore/BERTScore evaluations are completed with same resources as before.
@@ C.3 Resources
C.4 Choice for number of concepts $K$
Our choice of using $K=20$ concepts for all tokens was driven by the behaviour of reconstruction error of Semi-NMF on the training samples with different values of $K$, i.e. $||\mathbf{Z}-\mathbf{U}\mathbf{V}||_{2}^{2}$. We validate this behaviour on the four target tokens from main text in Fig. 11. We generally found $K=20$ as the minimal number of concepts where the reconstruction error drops by at least 50% from $K=0$.
We also conducted an ablation study to test how our method behaves with different values of number of concepts $K$. Fig. 12 presents the variation of test CLIPScore and Overlap score for $K\in\{10,20,30,50\}$ for two target tokens, ‘Dog’ and ‘Cat’. Our method can learn meaningful concepts for different values of $K$, evident by the consistently high CLIPScore. The Overlap score on the other hand tends to drop more consistently as the number of concepts increase but behaves stably for different choices. Nonetheless they indicate that our method can accommodate dictionaries of larger sizes without compromising the quality of learnt concepts, provided $K<<M$ (number of decomposed samples) and at the expense of greater user overhead.
Figure 11: Variation of reconstruction error with number of concepts $K$ for decompositions on different target tokens.
Figure 12: Test CLIPScore and Overlap score ablation with number of concepts $K$. CLIPScore remains consistently high and drops slightly only for very small $K$. Overlap score generally improves with higher $K$.
Appendix D Evaluation and extension to more tokens
We provide test data mean CLIPScore and BERTScore for top-1 activating concept for all baselines and more tokens: Baby, Car, Cake, and Bear in Tab. 9 (results in the main paper are reported for tokens Dog, Bus, Train, Cat in Tab. 1). Additionally, we also report the macro average over a set of 30 additional COCO-nouns apart from the 8 tokens with individually reported results, denoted as ‘Extra-30’. These nouns are single-token words with at least 40 predicted test samples. We put the filter of single-token words to keep consistency with the presented framework. Extension to multi-token words is straightforward but discussed separately in D.1. The lower bound criterion on test samples is to ensure average test CLIPScore is reliable for each target token. We only report CLIPScore for ‘Extra-30’ tokens as BERTScore evaluation was more expensive to conduct on large number of dictionaries.
We observe that we consistently obtain higher scores across for Semi-NMF and K-Means. We also report the overlap score between grounded words in Tab. 10 to illustrate the superiority of our method over the simple baseline. As previously noted, we observe a high overlap between grounded words with KMeans and Simple baselines compared to Semi-NMF/PCA. A low overlap should be encouraged, as it indicates the discovery of diverse and disentangled concepts in the dictionary. Among all the methods, Semi-NMF provides the most balanced concept dictionaries which are both meaningful (high CLIPScore/BERTScore) and diverse (low overlap).
@@ Appendix D Evaluation and extension to more tokens
Extra-30 CS top-1 ($\uparrow$) 0.516 $\pm$ 0.04 0.521 $\pm$ 0.03 0.626 $\pm$ 0.06 0.547 $\pm$ 0.06 0.637 $\pm$ 0.06 0.631 $\pm$ 0.06 0.763 $\pm$ 0.05
Table 9: Test data mean CLIPScore and BERTScore for top-1 activating concept CLIPScore denoted as CS, BERTScore denoted as BS for all concept extraction baselines considered. Analysis for layer $L=31$. Best score indicated in bold and second best is underlined. ‘Extra-30’ denotes a set of 30 additional single-token COCO nouns (apart from previous 8) with at least 40 predicted test samples by $f_{LM}$. For ‘Extra-30’ tokens we report the macro average and standard deviation of mean test data CLIPScore, taken over the set of 30 tokens.
@@
Extra-30 0.443 0.050 0.452 0.156
Table 10: Overlap/entanglement between grounded words of learnt concepts for different dictionary learning methods. Results are for additional tokens. Lower is better. Best score indicated in bold and second best is underlined. For ‘Extra-30’ tokens (additional 30 single-token COCO nouns) we report the macro average of Overlap score, taken over the set.
Statistical significance
The statistical significance of Semi-NMF w.r.t all other baselines and variants, for CLIPScore/BERTScore evaluation to understand representations of test samples is given in Tab/ 11 (for all tokens separately). We report the $p$-values for an independent two sided T-test with null hypothesis that mean performance is the same between Semi-NMF and the respective system. The results for Semi-NMF are almost always significant compared to Rnd-Words, Noise-Imgs, PCA. However for these metrics, Simple baseline, K-Means and Semi-NMF all perform competitively and better than other systems. Within these three systems the significance depends on the target token, but are often not significant in many cases.
@@
D.1 Extending to multi-token words
The presentation of our approach assumes that our token of interest $t$ is a single token. This poses no issues for words which are represented as single tokens but can raise some questions when we wish to extract concept representations for multi-token words. Our approach however, can be easily adapted to this setting. In particular, we extract representation of last token from first prediction of the multi-token sequence. Note that when filtering the training data for samples where ground-truth caption contains the token of interest, we now search for the complete multi-token sequence. The other aspects of the method remain unchanged. While there can also be other viable strategies, the rationale behind this adaptation is that the last token of our sequence of interest can also combines representations from previous tokens in the sequence. We add below results for such examples in Tab. 12 and 13. We observe behaviour consistent with the previous results with Semi-NMF extracting concept dictionaries with high CLIPScore and low overlap.
Multi-token word Rnd-Words Noise-Imgs Simple Semi-NMF GT-captions
Traffic light 0.516 $\pm$ 0.03 0.525 $\pm$ 0.03 0.664 $\pm$ 0.06 0.634 $\pm$ 0.05 0.744 $\pm$ 0.04
Cell phone 0.542 $\pm$ 0.04 0.547 $\pm$ 0.03 0.598 $\pm$ 0.04 0.598 $\pm$ 0.05 0.765 $\pm$ 0.06
Stop sign 0.533 $\pm$ 0.03 0.549 $\pm$ 0.03 0.617 $\pm$ 0.08 0.616 $\pm$ 0.05 0.775 $\pm$ 0.04
Table 12: Test mean CLIPScore ($\uparrow$) reported for top-1 activating concept for multi-token words. Higher scores are better. Best score in bold, second best is underlined.
Multi-token word Simple PCA K-Means Semi-NMF
Traffic light 0.704 0.050 0.579 0.174
Cell phone 0.623 0.051 0.746 0.164
Stop sign 0.461 0.058 0.704 0.109
Table 13: Overlap score ($\downarrow$) reported for top-1 activating concept for multi-token words. Higher scores are better. Best score in bold, second best is underlined.
Appendix E Additional visualizations
E.1 Concept grounding
The visual/textual grounding for all tokens in Tab. 1 are given in Figs. 13 (‘Dog’), 14 (‘Cat’), 15 (‘Bus’), 16 (‘Train’). All the results extract $K=20$ concepts from layer $L=31$. Similar to our analysis for token ‘Dog’ in main paper, for a variety of target tokens our method extracts diverse and multimodally coherent concepts encoding various aspects related to the token.
@@ E.1 Concept grounding
E.2 Local interpretations
Here, we qualitatively analyze the local interpretations of various decomposition methods, including PCA, k-means, semi-NMF, and the simple baseline strategy. We select these four as they produce coherent grounding compared to Rnd-Words and Noise-Img baselines. We decompose test sample representations on our learnt dictionary and visualize the top three activating components. Note that in the case of KMeans and Simple baseline, the projection maps a given test representation to a single element of the concept dictionary, the one closest to it. However, for uniformity we show the three most closest concept vectors for both. Figs. 17, 18, 19, 20, 21 are dedicated to interpretations of a single sample each, for all four concept extraction methods.
@@ E.2 Local interpretations
Appendix F Qualitative analysis for different layers
@@
Appendix G Analysis for visual tokens
@@ Appendix H Limitations
  • The current study is conducted mainly on visual captioning models. While we expect the key ideas to generalize to many other types of large multimodal models, this application of our approach to other types of LMMs remains to be explored/confirmed.
@@
Appendix I Broader societal impact